OpenStudy (anonymous):

Integrate. 1/sqrt(x^2+16)

OpenStudy (ash2326):

We have \[\int \frac{1}{\sqrt{x^2+16}} dx\] Do you know integration by substitution?

OpenStudy (anonymous):

yeah, I think did something wrong because I got \[1/4\ln \left| \sqrt{x^2+16} =x\right|\]

OpenStudy (ash2326):

Could you share your work here?

OpenStudy (anonymous):

I think I know why it's raised to the 1/4 power right. I think I get it. because if everything it raised to that the numbers that are over 4 go away.

OpenStudy (anonymous):

you did u sub? did you learn trig subs?

OpenStudy (anonymous):

Yes I did, I think I know my error.

OpenStudy (ash2326):

Could you try again? pr you want me to guide you? @CatLove9

OpenStudy (ash2326):

*or

OpenStudy (anonymous):

I got it, I forgot about sec^2

OpenStudy (ash2326):

@CatLove9 so what did you get this time?

OpenStudy (anonymous):

ln|sqrt(x^2+16)+x|

OpenStudy (ash2326):

there is a 1/4 missing in the integral. Would you check again?

OpenStudy (anonymous):

No there isn't I checked the answer in the book it's right because the dtheta is 4sec^2

OpenStudy (ash2326):

got it, sorry :(

OpenStudy (anonymous):

It's okay